# Milky Way Globular Clusters: Nurseries for Dynamically-Formed Binary Black Holes

### v1

@@ 3.3 Properties of Galaxies and Globular Clusters hosting Massive Binary Black Hole Formation
3.4 Redshift Evolution of BBH Merger Rates
Figure 4(a) illustrates the merger rate density of detectable dynamically-formed BBHs as a function of redshift. These systems are categorized by the mass of primary BHs: lighter BBHs ($_{1}\leq 100 M$) are shown in red solid lines, while systems with 100100 $_{\odot}< M$$_{1}\leq 1000$ M and $_{1}>1000 M$ are displayed in green and violet solid lines, respectively. The left panel presents results obtained using RAPSTER, while the right panel shows predictions from FASTCLUSTER.
The two CPS codes predict a slightly different number of detectable merging BBHs, as already shown in Fig. 2 and Table 1. Specifically, RAPSTER yields a lower merger rate estimate, characterized by a prominent peak at redshift of z $\sim 0.5$ followed by an exponential decline towards z = 0. In contrast, the FASTCLUSTER merger distribution peaks at z = 1, with a value slightly higher than the RAPSTER peak. Declines in both plots are attributed to the gradual decrease in the formation rate of GCs over time.
# BBH Mergers # Obs. BBH Mergers # Obs. BBH Mergers in the MW # IMBHs
RAPSTER 1.14 $10^{6}$ [2.39 $10^{4}$] 4.09 $10^{5}$ [5.44 $10^{3}$] 3.87 $10^{5}$ [5.13 $10^{3}$] 238 (4)
FASTCLUSTER 7.06 $10^{5}$ [1.21 $10^{3}$] 4.21 $10^{5}$ [18] 3.17 $10^{5}$ [15] 0
Table 1: Statistics of merging BBHs ($_{1}\leq 100 M$) from the GAMESH simulation. Columns indicate the total number of merging BBHs, the number of systems observable by terrestrial interferometers, the number of events formed within the MW, and the number of IMBHs ($_{1}>1000 M$) in the MW disk. The number of MBBHs ($100$ $_{\odot}< M$$_{1}\leq 1000$ M) is shown in square brackets, while values for IMBHs that migrated into the MW’s halo are in round parentheses.
RAPSTER predicts that merging MBBHs (with 100100 $_{\odot}< M$$_{1}\leq 1000$ M) are detectable at redshift z ¡ 2. In contrast, FASTCLUSTER predicts very few detectable MBBHs, and thus their corresponding rate density is not visible on the plot ($10^{-1}$ yr-1 cGpc-3).
@@ 3.4 Redshift Evolution of BBH Merger Rates
Interestingly, both CPS models show that the distribution of BBHs in the mass range of 100100 $_{\odot}< M$$_{1}\leq 1000$ M peaks at z $\sim 8$. In this mass range, RAPSTER consistently predicts a higher rate density for these systems.
An evident difference between the two CPS codes is that RAPSTER predicts the presence of IMBHs with masses greater than 1000 M that remain outside the current detectability window. The existence of these massive systems, particularly at high redshift, is significant because they could be the seed of supermassive BHs at the centers of galaxies. However, FASTCLUSTER does not predict the formation of these type of compact objects at all, though we are adopting the same initial conditions for both CPS codes. So, the predictions for the formation and merger rates of these systems are highly dependent on the adopted CPS model. A key parameter that might explain this difference is the high initial binary fraction. This parameter has a significant impact on RAPSTER predictions, particularly regarding the formation of IMBHs, while its influence is less pronounced in FASTCLUSTER. A full investigation into how the initial binary fraction affects IMBH formation will be addressed in future works to completely clarify its role.
Figure 4(b) (where we use the same color and line style conventions as in Fig. 4(a)) illustrates the distribution of the birth rate density for dynamically-formed BBHs. A direct comparison between the birth and merger rate density distributions provides significant insights into the dynamics of these binaries. Lighter BBHs ($_{1}<100 M$) consistently represent the majority of the BH population across all redshift values. Perhaps unexpectedly, some detectable BBHs originate at very high redshifts ($2\leq$ z $\leq 10$). This suggests that, according to both CPS models, the probability of observing systems formed in the early Universe is not negligible.
The strong similarity in the birth and merger rate trends for MBBHs ($100$ $_{\odot}< M$$_{1}\leq 1000$ M) and IMBHs (M >1000>1000 M) indicates an exceptionally rapid formation and coalescence timescale for these dynamically-formed systems. The formation of IMBHs is evident at redshifts above z $\sim 1.5$ for RAPSTER, whereas no IMBH formation is predicted at any redshift from FASTCLUSTER.
@@ 4 Discussion
The majority of existing works on GC formation focus on cosmological volumes (Pfeffer+2019; Ma+2020; Keller+2020; Doppel+2023; Bruel+2024), rather than on over-dense regions like our LG. As expected, our predictions for GC formation and abundance exceed those from cosmological simulations (DeLucia+2024; Bruel+2024). This is also seen in our GC formation rate, which is higher than the value inferred from the GWTC-3 catalog by a factor of a few tens (Fishbach+2023). This result highlights that our predictions should be considered as an upper limit when compared to values from other cosmological simulations and observations, which are limited by factors like mass resolution and current detector sensitivity, respectively.
@@
Regarding the age-metallicity relationship, our results show a marked difference from some recent studies. Kruijssen+2019; Kruijssen+2020 and DeLucia+2024, using the E-MOSAICS and Millennium simulations, respectively, find that their simulated GCs are systematically younger than the observed MW GCs, with an offset of $\sim 0.75$ Gyr. In contrast, our model shows good agreement with the observed age distribution of GCs. However, we find significant discrepancies in the metallicity distribution. Our results on this matter align with those of DeLucia+2024, who concluded that the metallicity distribution of simulated GCs is highly sensitive to the treatment of chemical enrichment. This discrepancy in metallicity may stem from our model’s lack of gas hydrodynamics and the resulting inability to accurately capture metallicity gradients within individual galaxies. Although our model reproduces the global properties of the MW, the discrepancy in the GC metallicity distribution suggests that our prescriptions for how metals are diluted and recycled within the ISM need to be revised for future GC formation studies. Specifically, our model may not be adequately capturing the processes that enrich the gas from which GCs form, especially in the early Universe, where most of the surviving GCs were born.
As shown in Section 3.4, our two CPS codes yield different results for the BBH merger rate density ($\mathcal{R}_{BBH}$). At the LVK Collaboration’s fiducial redshift of z = 0.2, we find merger rates of $\mathcal{R}_{BBH}\sim 21.2$ and 12.9 Gpc-3 yr-1 for RAPSTER and FASTCLUSTER, respectively. Both of these values slightly differ from the inferred value, which is in the range $22.5\div 37.5$ Gpc-3 yr-1 (GWTC4v2+2025). In any case, BBH mergers from this formation channel could be a potentially significant contribution to the total number of observed BH mergers. The difference in the merger rates between the two codes also highlights how sensitive the local merger rate density is to the specific GC evolution models.
@@ 5 Conclusions
  • Different CPS codes can lead to very different mass ranges for dynamically-formed BHs. RAPSTER predicts a mass range spanning from 4 to $\sim$ 4000 M, while FASTCLUSTER constrains the mass range to a narrower window of $15\div 200$ M. This disparity underscores the heavy dependence of model outcomes on the internal physics of the CPS code.
@@
Future work will tackle the problem of MBBH formation by using a more sophisticated cosmological simulation, such as dustyGadget (Graziani+2020c), which may help to reconcile the predictions of our cluster formation model with observations.
@@ Acknowledgements.
F.A. acknowledges support from Sapienza and Tor Vergata University of Rome. F.A. also acknowledges support from a MSCA-GRU action funded by the European Union during the visiting period (2024 October–November) at Johns Hopkins University, Baltimore (USA). This work was supported by the Piano Nazionale di Ripresa e Resilienza (PNRR). We are grateful to Angela Adamo for her insightful suggestions on globular cluster modeling. E.B. and K.K. are supported by NSF Grants No. AST-2307146, PHY-2513337, PHY-090003, and PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, E.B.], by the Simons Foundation International [SFI-MPS-BH-00012593-02], and by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. K.K. is supported by the Onassis Foundation - Scholarship ID: F ZT 041-1/2023-2024. We have benefited from the publicly available programming language Python, including the numpy, matplotlib, and scipy packages.
Appendix A Survived GC Mass-Metallicity Distribution
@@
Figure 5 illustrates the mass-metallicity distribution of GCs that survive to redshift z = 0. The number and final masses of these GCs are in good agreement with observational data, as shown by the significant overlap between the simulated and observed GC mass distributions in the top histogram. The majority of surviving GCs experience substantial mass loss, typically around one order of magnitude, given their initial masses exceeding $10^{5}$ M. While our sample generally covers the observed mass range, we note some difficulty in populating the extreme ends of the mass spectrum. Specifically, the under-representation of the highest masses may imply a need for more massive initial clusters in our seeding population. Conversely, for the less massive GCs, it is likely that these samples originated from initial masses below $10^{5}$ M, thereby falling outside the GC mass range considered in our study. The filled points in Fig. 5 are color-coded by the initial central stellar number density of these GCs, spanning a wide range of values ($10^{5}\div 10^{7}$ pc-3). This broad distribution confirms that our selection process of surviving GCs introduces no discernible bias towards a limited range of initial densities. Interestingly, despite the presence of highly compact GCs, we find that only four IMBHs form in the MW halo. This finding suggests that high compactness alone is insufficient for IMBH formation; a sufficiently high initial cluster mass (around $10^{6}$ M) is also a necessary condition (see Section 3.3). Higher masses are crucial for refueling hierarchical BH merger chains, while higher compactness promotes BH encounters and accelerates IMBH formation. The GCs capable to produce IMBHs have been marked with stars in the plot.
@@
Figure 5: Mass-metallicity distribution of surviving MW GCs. The color bar indicates the central stellar number density of GCs. The red dashed line in the top histogram highlights the mass distribution of the observed GCs in our MW (Kruijssen+2015).

### v2

@@ 3.3 Properties of galaxies and globular clusters hosting massive binary black hole formation
3.4 Redshift evolution of binary black hole merger rates
Figure 4(a) illustrates the merger rate density of detectable dynamically formed BBHs as a function of redshift. These systems are categorized by the mass of primary BHs: lighter BBHs ($_{1}\leq 100 M$) are shown in red solid lines, while systems with $100$ $_{\odot}< M$$_{1}\leq 1000$ M and $_{1}>1000 M$ are displayed in green and violet solid lines, respectively. The left panel presents results obtained using RAPSTER, while the right panel shows predictions from FASTCLUSTER.
Table 1: Statistics of merging BBHs and IMBHs from the GAMESH simulation. [^3] Columns indicate the total number of merging BBHs ($\rm M_{1}\leq 100\,\rm M_{\odot}$), the number of systems observable by terrestrial interferometers, the number of events formed within the MW, and the number of IMBHs ($\rm M_{1}>1000\,\rm M_{\odot}$) in the MW disk. The number of MBBHs ($100$ $_{\odot}< M$$_{1}\leq 1000$ M) is shown in square brackets, while values for IMBHs that migrated into the MW’s halo are in round parentheses.
Model # BBH Mergers # Obs. BBH Mergers # Obs. BBH Mergers MW # IMBHs
RAPSTER $1.14\times 10^{6}$ $2.39\times 10^{4}$] $4.09\times 10^{5}$ $5.44\times 10^{3}$] $3.87\times 10^{5}$ $5.13\times 10^{3}$] 238 (4)
FASTCLUSTER $7.06\times 10^{5}$ $1.21\times 10^{3}$] $4.21\times 10^{5}$ [18] $3.17\times 10^{5}$ [15] 0
The two CPS codes predict a slightly different number of detectable merging BBHs, as already shown in Fig. 2 and Table 1. Specifically, RAPSTER yields a lower merger rate estimate, characterized by a prominent peak at redshift of z $\sim 0.5$ followed by an exponential decline toward z = 0. In contrast, the FASTCLUSTER merger distribution peaks at z = 1, with a value slightly higher than the RAPSTER peak. Declines in both plots are attributed to the gradual decrease in the formation rate of GCs over time.
RAPSTER predicts that merging MBBHs (with $100$ $_{\odot}< M$$_{1}\leq 1000$ M) are detectable at redshift z ¡ 2. In contrast, FASTCLUSTER predicts very few detectable MBBHs, and thus their corresponding rate density is not visible on the plot (¡ $10^{-1}$ yr-1 cGpc-3).
@@ 3.4 Redshift evolution of binary black hole merger rates
Interestingly, both CPS models show that the distribution of BBHs in the mass range of $100$ $_{\odot}< M$$_{1}\leq 1000$ M peaks at z $\sim 8$. In this mass range, RAPSTER consistently predicts a higher rate density for these systems.
An evident difference between the two CPS codes is that RAPSTER predicts the presence of IMBHs with masses greater than 1000 M that remain outside the current detectability window. The existence of these massive systems, particularly at high redshift, is significant because they could be the seed of supermassive BHs at the centers of galaxies. However, FASTCLUSTER does not predict the formation of these type of compact objects at all, though we are adopting the same initial conditions for both CPS codes. So, the predictions for the formation and merger rates of these systems are highly dependent on the adopted CPS model. The discrepancy in the formation of IMBHs between the two codes stems from their fundamental treatment of cluster dynamics. FASTCLUSTER operates as a semi-analytical framework where cluster properties (such as mass M and half-mass radius rh) evolve through average differential equations (e.g., Gieles et al. 2021). This approach lacks the resolution to capture the stochastic, runaway growth of a single compact object. Consequently, the absence of IMBHs in FASTCLUSTER is not a numerical error, but a direct consequence of its mean-field statistical nature, as it is not designed to capture those rare cases where a single object rapidly grows within a tightly packed cluster.
A key factor driving the discrepancy is the sensitivity of the results to the initial binary fraction fb. FASTCLUSTER global evolution remains largely unaffected by variations in fb, whereas the formation of IMBHs in RAPSTER is strictly dependent on it. This behavior highlights the different physical treatments: RAPSTER explicitly models binary-mediated heating. In a high-fb environment, the continuous hardening of BH binaries provides the necessary kinetic energy to support the BH subsystem against rapid contraction, maintaining a high-density ”active” core where hierarchical mergers can proceed.
The fact that IMBH formation in RAPSTER vanishes at lower fb confirms that the runaway growth we report is a dynamically driven process fueled by cumulative binary interactions effects that are simplified in FASTCLUSTER framework, which treats binary interactions as isolated events without collective dynamical feedback. A full investigation into how the initial binary fraction affects IMBH formation will be addressed in future works to completely clarify its role.
Figure 4(b) (where we use the same color and line style conventions as in Fig. 4(a)) illustrates the distribution of the birth rate density for dynamically formed BBHs. A direct comparison between the birth and merger rate density distributions provides significant insights into the dynamics of these binaries. Lighter BBHs ($_{1}<100 M$) consistently represent the majority of the BH population across all redshift values. Perhaps unexpectedly, some detectable BBHs originate at very high redshifts ($2\leq$ z $\leq 10$). This suggests that, according to both CPS models, the probability of observing systems formed in the early Universe is not negligible.
The strong similarity in the birth and merger rate trends for MBBHs ($100$ $_{\odot}< M$$_{1}\leq 1000$ M) and IMBHs (M >1000>1000 M) indicates an exceptionally rapid formation and coalescence timescale for these dynamically formed systems. The formation of IMBHs is evident at redshifts above z $\sim 1.5$ for RAPSTER, whereas no IMBH formation is predicted at any redshift from FASTCLUSTER.
@@ 4 Discussion
The majority of existing works on GC formation focus on cosmological volumes (Pfeffer et al. 2019; Ma et al. 2020; Keller et al. 2020; Doppel et al. 2023; Bruel et al. 2024), rather than on over-dense regions similar to our LG. As expected, our predictions for GC formation and abundance exceed those from cosmological simulations (De Lucia et al. 2024; Bruel et al. 2024). This is also seen in our GC formation rate, which is higher than the value inferred from the GWTC-3 catalog by a factor of a few tens (Fishbach and Fragione 2023). This result highlights that our predictions should be considered as an upper limit when compared to values from other cosmological simulations and observations, which are limited by factors such as mass resolution and current detector sensitivity, respectively.
@@
Regarding the age-metallicity relationship, our results show a marked difference from some recent studies. Kruijssen et al. (2019, 2020) and De Lucia et al. (2024), using the E-MOSAICS and Millennium simulations, respectively, find that their simulated GCs are systematically younger than the observed MW GCs, with an offset of $\sim 0.75$ Gyr. In contrast, our model shows good agreement with the observed age distribution of GCs. However, we find significant discrepancies in the metallicity distribution. Our results on this matter align with those of De Lucia et al. (2024), who concluded that the metallicity distribution of simulated GCs is highly sensitive to the treatment of chemical enrichment. This discrepancy in metallicity may stem from our model’s lack of gas hydrodynamics and the resulting inability to accurately capture metallicity gradients within individual galaxies. Although our model reproduces the global properties of the MW, the discrepancy in the GC metallicity distribution suggests that our prescriptions for how metals are diluted and recycled within the ISM need to be revised for future GC formation studies. Specifically, our model may not be adequately capturing the processes that enrich the gas from which GCs form, especially in the early Universe, where most of the surviving GCs were born.
Stellar mergers in young clusters (e.g., Di Carlo et al. 2019) can produce gap-BHs by creating massive stars that avoid the PI limit. However, our results suggest that in GCs the formation of objects exceeding 50 M is strictly tied to the cluster’s structural evolution, specifically requiring a central stellar density $10^{7}$ pc-3 and a total mass $10^{6}$ M. This highlights a higher degree of ”environmental specialization” compared to young clusters.
[…]
As shown in Section 3.4, our two CPS codes yield different results for the BBH merger rate density ($\mathcal{R}_{BBH}$). At the LVK Collaboration’s fiducial redshift of z = 0.2, we find merger rates of $\mathcal{R}_{BBH}\sim 21.2$ and 12.9 Gpc-3 yr-1 for RAPSTER and FASTCLUSTER, respectively. Both of these values slightly differ from the inferred value, which is in the range 22.537.522.5-37.5 Gpc-3 yr-1 (The LIGO Scientific Collaboration et al. 2025a). In any case, BBH mergers from this formation channel could be a potentially significant contribution to the total number of observed BH mergers. The difference in the merger rates between the two codes also highlights how sensitive the local merger rate density is to the specific GC evolution models.
@@ 5 Conclusions
  • Different CPS codes can lead to very different mass ranges for dynamically formed BHs. RAPSTER predicts a mass range spanning from 4 to $\sim$ 4000 M, while FASTCLUSTER constrains the mass range to a narrower window of 1520015-200 M. This disparity underscores the heavy dependence of model outcomes on the internal physics of the CPS code.
@@
Future work will tackle the problem of MBBH formation by using a more sophisticated cosmological simulation, such as dustyGadget (Graziani et al. 2020b), which may help to reconcile the predictions of our cluster formation model with observations.
@@ Acknowledgements.
F.A. acknowledges support from Sapienza and Tor Vergata University of Rome. F.A. also acknowledges support from a MSCA-GRU action funded by the European Union during the visiting period (2024 October–November) at Johns Hopkins University, Baltimore (USA). This work was supported by the Piano Nazionale di Ripresa e Resilienza (PNRR). We are grateful to Angela Adamo for her insightful suggestions on globular cluster modeling. E.B. and K.K. are supported by NSF Grants No. AST-2307146, PHY-2513337, PHY-090003, and PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, E.B.], by the Simons Foundation International [SFI-MPS-BH-00012593-02], and by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. K.K. is supported by the Onassis Foundation - Scholarship ID: F ZT 041-1/2023-2024. MM and ST acknowledge financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017 and from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. ST acknowledges financial support from the Alexander von Humboldt Foundation for the Humboldt Research Fellowship. We have benefited from the publicly available programming language Python, including the numpy, matplotlib, and scipy packages.
Appendix A Survived globular cluster mass-metallicity distribution
@@
Figure 5 illustrates the mass-metallicity distribution of GCs that survive to redshift z = 0. The number and final masses of these GCs are in good agreement with observational data, as shown by the significant overlap between the simulated and observed GC mass distributions in the top histogram. The majority of surviving GCs experience substantial mass loss, typically around one order of magnitude, given their initial masses exceeding $10^{5}$ M. While our sample generally covers the observed mass range, we note some difficulty in populating the extreme ends of the mass spectrum. Specifically, the under-representation of the highest masses may imply a need for more massive initial clusters in our seeding population. Conversely, for the less massive GCs, it is likely that these samples originated from initial masses below $10^{5}$ M, thereby falling outside the GC mass range considered in our study. The filled points in Fig. 5 are color-coded by the initial central stellar number density of these GCs, spanning a wide range of values ($10^{5}-10^{7}$ pc-3). This broad distribution confirms that our selection process of surviving GCs introduces no discernible bias toward a limited range of initial densities. Interestingly, despite the presence of highly compact GCs, we find that only four IMBHs form in the MW halo. This finding suggests that high compactness alone is insufficient for IMBH formation; a sufficiently high initial cluster mass (around $10^{6}$ M) is also a necessary condition (see Section 3.3). Higher masses are crucial for refueling hierarchical BH merger chains, while higher compactness promotes BH encounters and accelerates IMBH formation. The GCs hosting the formation of IMBHs have been marked with stars in the plot.
@@
Figure 5: Mass-metallicity distribution of surviving MW GCs. The color bar indicates the initial central stellar number density of GCs, while star-shaped symbols mark the GCs hosting an IMBH. The dashed red line in the top histogram highlights the mass distribution of the observed GCs in our MW (Kruijssen 2015).
Appendix B Sensitivity of globular cluster lifetimes to orbital assumptions
As shown in Fig. 6, the cluster lifetimes are sensitive to their spatial distribution within the galaxy. In our reference model (black histogram), we assume that all GCs lie at an orbital radius of rg = rh (see section 2.2). To quantify the impact of this assumption, we performed a sensitivity test at z=6 by varying the orbital radii of $\pm$30%.
In the rg = 0.7 rh case (red histogram), the increased frequency of encounters with molecular clouds in the denser inner regions leads to significantly shorter lifetimes, with the distribution peak shifting below 100 Myr. Conversely, at rg = 1.3 rh (green histogram), the reduced environmental density allows for much longer survival times, often exceeding 1 Gyr.
In Fig. 6, we also show the average BBH delay time ($\sim$ 400 Myr, i.e., the time elapsed between the formation of the BBH and its merger). This value aligns almost perfectly with the peak of the lifetime distribution in our reference model. In contrast, the rg = 0.7 rh scenario would imply that a large fraction of GCs are disrupted before their internal BBH populations have sufficient time to merge, potentially underestimating the GW event rate. In the rg = 1.3 rh case, the extended cluster lifetimes would allow for a higher number of dynamical interactions leading to a higher BBH merger rate. These results indicate that while rg = rh works as a representative average environment.
Figure 6: Distributions of GC lifetimes as a function of their orbital distance from the galactic center. Black line represents our fiducial model (rg = rh), while red and green histograms indicate deeper orbits (rg = 0.7 rh) and more peripheral orbits (rg = 1.3 rh), respectively. Vertical dashed red line indicates the BBH mean delay time.
Appendix C Impact of the gas surface density threshold on the globular cluster formation environments
(a) Without threshold
(b) $\Sigma_{g}$ ¿ 100 M pc-2
(c) $\Sigma_{g}$ ¿ 5000 M pc-2
(d) $\Sigma_{g}$ ¿ $10^{4}$ M pc-2
Figure 7: Age-metallicity distribution of all potential GC formation environments within our LG-like volume, assuming different values for the $\Sigma_{g}$ threshold: (a) without threshold, (b) $\Sigma_{g}$ ¿ 100 M pc-2, (c) $\Sigma_{g}$ ¿ 5000 M pc-2, (d) $\Sigma_{g}$ ¿ $10^{4}$ M pc-2). Red dots with error bars correspond to individual GCs observed in our Galaxy (Kruijssen 2015).
The gas surface density threshold is a critical parameter in distinguishing the galaxies of our LG-like volume capable to form GCs. As illustrated in Fig. C, varying this threshold significantly impacts the number of host environments. We tested four values - above and below our main configuration - to evaluate the sensitivity of our results to this parameter.
If the threshold is high (>1000>1000 M pc-2), GC formation is restricted only to massive galaxies like the MW and Andromeda, failing to account for the spread of observational data. Indeed, we could not explain the metal-poor GCs. Conversely, an excessively low (or absent) threshold leads to overpredict the specific frequency of GCs in the LG, since also extremely metal-poor environments (Z $\approx$ 10-4 Z) become GC hosting galaxies. These variations also inevitably influence the GC survival fraction within the MW halo. In the cases with $\Sigma_{g}>5000$ M pc-2, GC formation is suppressed in almost all progenitor galaxies except for the main MW progenitor. Consequently, we found $\sim 130$ surviving GCs in the MW halo at z = 0 originate exclusively from the MW disk. This scenario fails to account for the significant population of ex situ surviving GCs known to be accreted from satellite galaxies. Conversely, the other two cases allow even small satellite galaxies to form GCs. This leads to an overproduction ($\sim 300$ surviving GCs in the MW halo), as every minor and major merger event efficiently injects GCs into the halo.
Our reference model, with $\Sigma_{g}>1000$ M pc-2, yields $185$ surviving GCs in the MW halo, providing the best agreement with both observed GCs and the expected contribution from accreted satellites. We therefore selected this value as our fiducial configuration.